Suppose that business travelers and vacationers have the following demand for airline tickets from Karachi to Dubai:

 

Price

Quantity Demanded

(Business Travelers)

Quantity Demanded

(vacationers)

$150

2100 Tickets

1000 Tickets

$200

2000

 800

$250

$300

1900

1800

 600

 400

 

 

 

 

 

a.      As the price of tickets rises from $150 to $200, what is the price elasticity of demand for (i) business travelers and (ii) vacationers? (Use the midpoint method in your calculations.)

b.      Why might vacationers have a different elasticity from business travelers?

c.      What would you recommend based on the above information for airline to increase the revenue?


1
Expert's answer
2021-05-18T19:35:13-0400

"Soln,"

a. Price elasticity of demand (midpoint method)

"=\frac {Q2-Q1}{P2-P1}\times(\frac {P2+P1}{2}\times\frac{2} {Q2+Q1})"


"=\frac {Q2-Q1}{P2-P1}\times\frac{P2+P1}{Q1+Q2}"


For business travellers, When; P1=$150 ,Q1=2100 Tickets.

when; P2= $200 ,Q2=2000 Tickets.

i. Price elasticity for business travellers."=\frac{2000-2100}{200-150}\times\frac{150+200}{2100+2000}"


"=\frac{-100}{50}\times\frac{350}{4100}=-0.17"


"=0.17"


ii. For vacationers, When; P1=$150 ,Q1=1000 Tickets.

When, P2=$200 ,Q2=800 Tickets.


"=\frac {Q2-Q1}{P2-P1}\times \frac{P2+P1}{Q1+Q2}"


"=\frac {800-1000}{200-150}\times\frac{150+200}{1000+800}=\frac{-200}{50}\times\frac{350}{1800}=-0.78"


"=0.78"

b. Vacationers have a higher elasticity of demand for air tickets as compared to the business travellers. For vacationers, travelling is less of a necessity compared to business travellers ,hence they have a longer time frame for travel while business travellers have very little time to be in between places. Moreover vacationers have several other effective and efficient travel substitutes in the form of road transport and Rail.


c. The airline when keen to increase their revenues should give priority travel to business travellers since at whatever prices offered they buy more travel tickets than the vacationers.

Also the airline should segment vacation and business travel flights and strive to sell these separate identities to travellers looking for either of the extremes.